OpenStudy (anonymous):

Let f(x)=5x-6 and g(x)=7x-2. find f(x)+g(x) eeek how do i solve this?

OpenStudy (***[isuru]***):

f(x) = 5x-6 ----(1) g(x)=7x-2 ----(2) (1) +(2) f(x) + g(x) = 5x -6 + 7x -2.. can u continue ?

jhonyy9 (jhonyy9):

courage @memirobb !

jhonyy9 (jhonyy9):

just assum the same terms

OpenStudy (***[isuru]***):

u can add and subtract the like terms... i mean the terms which include "x" for an example if there is something like this 8x + 3x u can add 8 and 3 together and write 11x if there is 8x - 3x u can sub 3 from 8 and write the answer as 5x in ur question if u bring up the like terms close together f(x) + g(x) = 5x -6 + 7x -2. write the terms with same variable close = 5x + 7x -6 - 2 now its like adding 8x and 3x like i mention above... u can do the same and simplify 5x + 7x.. -6 -2 ... i think u dont have a problem simplifying that part

jhonyy9 (jhonyy9):

@memirobb what is the problem ? are you here again ?

OpenStudy (anonymous):

okayy i gotcha.. is it 12x-4? @***[ISURU]***